# A regularized continuum model for traveling waves and dispersive shocks of the granular chain

### v1

@@ 2 Models and theoretical setup
In this work we focus on granular lattice dynamical systems. The non-dimensional equations of motion (where the constant elastic and geometric prefactors have been absorbed through a suitable rescaling) is given by the following differential-difference equations (DDEs) [5, 11, 26]:
@@
where $\delta_{0}$ denotes the precompression constant, $u_{n}$ the displacement of the $n$th particle from its equilibrium position, and $\left[f\right]_{+}=\text{max}(f,0)$, models the fact that there is no force when the particles come out of contact. The case of $p=3/2$ corresponds to a lattice of spherical particles. It is relevant to mention in passing that in other settings such as O-rings, cylindrical particles or hollow spheres the nature of the force (and hence exponent) may vary [27] and hence we will maintain $p$ as a general parameter in our considerations herein. Equation (2.1) possesses traveling wave solutions [5, 14, 11, 12], even though their exact form is not known analytically (but for relevant approximations see e.g., [28]). Moreover, numerical investigations of Eq. (2.1) reveal the formation of dispersive shocks in certain regimes [7, 8, 9, 26]. In a recent work, we studied the DSWs produced by Eq. (2.1) when $\delta_{0}\neq 0$, making use of two different approximations for it: the Toda lattice and the Korteweg-de Vries equation [25]. In the present work, we focus on the case where $\delta_{0}=0$, in which case the model does not admit a meaningful linear limit in the case of zero background.
It will be convenient to work with the strain variable $r_{n}=u_{n-1}-u_{n}$, which has a physically meaningful interpretation in terms of granular crystals (the amount of compression between adjacent particles). This allows one to rewrite Eq. (2.1) as
$\ddot{r}_{n}=\left(r_{n+1}\right)^{p}-2\left(r_{n}\right)^{p}+\left(r_{n-1}\right)^{p}.$ (2.2)
The linearized dispersion relation for small-amplitude wave solutions of the form $r_{n}(t)=A+Be^{i(kn-\omega t)}$, with $|B/A|\ll 1$, is given by
@@
$r_{TT}=\left(r^{p}\right)_{XX}+\frac{\epsilon^{2}}{12}\left(r^{p}\right)_{XXXX}=\left(r^{p}\right)_{XX}+\mathcal{O}\left(\epsilon^{2}\right),$ (2.10)
@@
$\left(r^{p}\right)_{XXXX}=\left(\left(r^{p}\right)_{XX}\right)_{XX}=\left(r_{TT}-\mathcal{O}\left(\epsilon^{2}\right)\right)_{XX}=r_{TTXX}+\mathcal{O}\left(\epsilon^{2}\right).$ (2.11)
@@ 3.1 Derivation of solitary waves
where denotes differentiation with respect to $Z$ throughout this section, and $a,b$ are two integration constants. When $p$ takes on half-integer values, it is then convenient to apply the transformation $R=g^{2}$ so that Eq. (3.2) becomes
@@
where $Z_{0}$ is a integration constant. Indeed, the relevant calculation arises in the process of identifying the bright solitary waves of the general power variant of the nonlinear Schrödinger equation [35]. Recall that since $R=g^{2}$, the corresponding traveling wave solution for $R$ assumes the form
@@ 3.2 Comparison of traveling waves of different models
Figure 2 showcases the numerical comparison of 4 traveling wave solutions (3 approximations and 1 “exact” —again, to a prescribed tolerance—) for the cases of $p=3/2$,$2$ $3$. Each of the traveling wave solutions which arise in periodic (cosinusoidal) form is plotted only for one period of the corresponding cosine function, beyond which the solution is taken to be zero. Note that, while the continuum models themselves depend on the parameter $\epsilon$, the corresponding predictions is independent of $\epsilon$ once returning to the original lattice variables, see e.g. Eq. (3.7). Thus, we do not need to “select” any particular value of $\epsilon$ when comparing solitary waves.
@@
For the values of $p$ considered here, the approximation from (3.7) is meaningfully proximal to the exact solution, as are the other models. See the black solid curve of Fig. 2. In terms of the error $E=\frac{1}{N}\sum_{n}^{N}|r_{n}-R(\epsilon n)|$, where $r_{n}$ is an “exact” traveling wave and $R$ a traveling wave from one of the PDE models, the approximation errors are of the same magnitude. For example, with $p=3/2$ the errors are $E=0.034$ for the strain variant of the model of Nesterenko (3.11), $E=0.030$ for the model of Pikovsky (2.9), and $E=0.046$ for the model of considered in this paper, namely the regularized model of (2.12). The errors for other values of $p$ are similar. Recall that the solitary wave solution of the granular chain has a double exponential decay in the tails [37, 38], whereas the approximation from (3.7) has only exponential decay. The approximations (3.9) and (3.13) have finite support. See Fig. 2(d-f), which shows the solutions in a semi-log scale where the decay rate can be better discerned. While the quantitative differences between the actual solution and approximation are similar for all cases, it still remains an interesting open question if, in terms of rigorous error bounds, any of the three considered here is the “most” accurate. It is worth mentioning that here we have only examined the stationary (in the co-traveling frame) aspect of solitary waves. Towards the end of our presentation, we will return to the dynamical properties of these models, as concerns the prototypical structure considered herein, namely the DSW.
@@
Figure 2: Comparison of the traveling solitary waves in different continuum models for different values of the parameter $p$: The leftmost, middle, and rightmost columns denote the cases of $p=3/2,2,3$, respectively. The blue lines (with circles) denote the “exact” solitary of Eq. (3.15), and the red dashed line, pink dotted line, black solid line refer to the solitary waves associated with the Nesterenko (2.9), Ahnert-Pikovsky (3.11), and the regularized continuum model (2.12), respectively. The first row shows the comparison of the solitary waves of different continuum models in their respective standard scale, while the second row depicts the semi-log scale of all solitary waves. Note that Pikovsky’s and Nesterenko’s approximations for the solitary waves are only plotted over the period of the respective cosine.
@@ 4.1 Case $p=3/2$
where, as before, primes denote differentiation with respect to $Z$. A direct integration of Eq. (4.1) yields,
@@
Note that this is only a two parameter family of periodic waves, and hence is not the most general form. In the derivation of this formula, one of the integration constants was set to zero, see Eq. (4.1). We were unable to find a analytical formula in the general, three parameter, case.
@@ 4.2 Case $p=2$
Note, in contast to the previous case, this ODE has three free parameter $(B,C,c)$. A direct integration of Eq. (4.4) yields
@@ 4.3 Case $p=3$
Clearly $\mu<0$, and then we assume the following order of the four roots,
@@
Here, we need to take the root $R_{4}^{+}$ and ignore $R_{4}^{-}$ to avoid the issue of negative soliton amplitude. Finally, substituting $R_{4}^{+}$ into Eq. (4.16) we obtain an explicit soliton amplitude formula for $p=3$,
@@ 5 Conservation laws
respectively. Moreover, at the strain level, we notice that the PDE (2.12) itself forms a conservation law,
@@ 6.1 Theoretical preliminaries
The idea of modulation theory is to consider slow modulations of the parameters that completely determine the periodic solutions and derive their associated governing equations. To this end, we first introduce the wavenumber $K$ and frequency $\Omega$ to rewrite the traveling-wave ansatz as $r(X,T)=R(\theta)$, with the phase variable $\theta=(KX-\Omega T)/\epsilon$ where $\Omega=c_{p}K$, to express our periodic solutions. In the case of $p=2$, the periodic solution in terms of $\theta$ is,
$R\left(\theta\right)=R_{2}+\left(R_{3}-R_{2}\right)\text{cn}^{2}\left(\frac{\sqrt{2\left(R_{3}-R_{1}\right)}}{cK}\left(\theta-\theta_{0}\right),m\right),$ (6.1)
@@
$2\pi=\int_{0}^{2\pi}\text{d}\theta=2\int_{R_{2}}^{R_{3}}\frac{dR}{R_{\theta}}=2\int_{R_{2}}^{R_{3}}\frac{dR}{\sqrt{\frac{8}{K^{2}c^{2}}\left(R_{1}-R\right)\left(R_{2}-R\right)\left(R_{3}-R\right)}}=\frac{2c_{p}KK_{m}}{\sqrt{2\left(R_{3}-R_{1}\right)}},$ (6.2)
@@
$R\left(\theta\right)=R_{2}+\left(R_{3}-R_{2}\right)\text{cn}^{2}\left(\frac{K_{m}}{\pi}\theta,m\right),$ (6.4)
@@ 6.2 Derivation of modulation equations
To derive the modulation equations, we will use the method of averaged Lagrangian [42]. We first note that the PDE model in its displacement form, see Eq. (5.3), can be obtained through a variational principle. In the case of $p=2$ the Lagrangian density $\mathbb{L}$ is
$\mathbb{L}=\frac{1}{2}\left(u_{T}\right)^{2}+\frac{\epsilon^{2}}{24}u_{XX}u_{TT}-\frac{\left(u_{X}\right)^{3}}{3}.$ (6.9)
@@
Note that, with this ansatz, all terms can be expressed in terms of the traveling wave at the strain level $R$. In paricular, $R=R(\theta)$ satisifies
$K^{2}R_{\theta}^{2}=-\frac{8}{c_{p}^{2}}R^{3}+12R^{2}-24BR-12C,$ (6.13)
@@
$\left(\frac{c_{p}}{12}\left(2W+c_{p}W_{c_{p}}\right)+\frac{2c_{p}}{K}\left(-\frac{1}{2}\beta^{2}+\frac{1}{2}C+\beta B\right)\right)_{T}+\left(\frac{c_{p}^{2}}{12}\left(W_{c_{p}}c_{p}+W\right)+\frac{2c_{p}^{2}}{K}\left(-\frac{1}{2}\beta^{2}+\frac{1}{2}C+\beta B\right)\right)_{X}$ $=0,$ (6.25a)
@@ 7 Riemann problems and DSW fitting
7.1 Numerical simulation and DSWs
We begin by establishing a connection between the two Riemann problems mentioned above and specify in detail what initial conditions are applied and what necessary conditions they have to fulfill. While the classic Riemann problem has the single step, as defined in Eq. (2.4), we are working with periodic boundary conditions, so we employ an initial condition that has a step-up and then a step-down. In particular, for the continuum PDE (7.4) we use the following “box-type” initial data for the strain variable $r$:
$r(X,0)=\begin{cases}r^{-}\,,&a<X<b\,,\\ r^{+}\,,&X<a~{}\vee~{}X>b\,,\end{cases}$ (7.1)
where $a,b\in\mathbb{R}$ are two real constants with $a<b$. In what follows, we assume that $r_{-}>r_{+}$, in which case we will obtain right-moving DSWs (the case of $r_{-}<r_{+}$ will yield left-moving DSWs). For the numerical approximation of the PDEs, we employ a smooth approximation of the above initial condition as
The initial data for the DDE simulation is chosen to be consistent with Eq. (7.2), which we discuss in detail later in Sec. 7.4. Figure 3 displays the simulation of the Riemann problem for the DDE. A waveform emerges at the right upper corner of the ”box-type” initial data. This waveform connects left state ($r^{-}$) to the right state ($r^{+}$) and expands as time increases. If one zooms into a small spatial-window of the waveform, it will appear as though the wave is periodic. The underlying parameters of the periodic wave will depend on where in the lattice you zoom, and hence the waveform is modulated. This is the so-called DSW. At the front of the wave (called the leading edge) the wave closely resembles a solitary wave, and it travels with a near constant speed $s^{+}$ and has some amplitude $a^{+}$. On the other hand, the back part of the DSW (called the trailing edge) is characterized by a linear wave with some wavenumber $k^{-}$ traveling at a speed $s^{-}$ with amplitude that vanishes as the trailing edge is approached from within the DSW. For any finite time snapshot of the DSW, there will be linear waves that are located around the background value of $r^{-}$. These linear waves are not a part of the DSW per se, and vanish for $t\rightarrow\infty$. Figure 3(a) shows a time snapshot of the DSW, where the key elements of the DSW, namely the leading and trailing edge can be seen, along with the linear waves that are always present for any finite time snap shot. Figure 3(b) shows an intensity plot of the spatio-temporal evolution where the expanding oscillatory region (i.e., the DSW) can be seen. The dashed lines in panel (b) are analytical approximations of the leading and trailing edges that stem from the so-called DSW fitting method. Before discussing those approximations, we discuss some preliminaries involving the the dispersionless system.
Figure 3: Numerical simulation of the Riemann problem: The left panel shows the profile of the DSW of the continuum model (2.12) at $T=50$, while the right panel depicts the density plot of the DSW, and the two black dashed lines in the density plot represent the leading (upper) and the trailing (lower) edges of the DSW, respectively. In particular, the upper and the lower dashed black lines depict $X=s^{+}T$ and $X=s^{-}T$ where $s^{+},s^{-}$ are obtained based on Eqs. (7.21). Meanwhile, notice that the parameters of $r^{+},r^{-}$ in the Riemann initial condition (7.1) are set to be $r^{+}=0.8$, and $r^{-}=1$ in the numerical simulation.
@@ 7.2 Riemann invariants of the dispersionless system
We recall and notice that the continuum model Eq. (2.12) is a bi-directional model. Therefore, to numerically simulate the associated Riemann problem, we not only need the initial data of the strain variable $r$, which is given in Eq. (7.1), but we also need an initial condition for the density variable $\rho$ which will be connected to the time derivative of the strain in the upcoming discussion. As a preliminary step to understand how the initial conditions are properly set up, we first need to understand the following dispersionless averaged version of the continuum model (2.12) where [22] claims that the Whitham modulation equations at both harmonic and solitonic edges should agree with, namely they will both read
@@
The dispersionless first-order system simply stems from the following formulation of our PDE model,
$\left(r-\frac{\epsilon^{2}}{12}r_{XX}\right)_{T}-\rho_{X}$ $=0,$ (7.4)
$\rho_{T}$ $=\left(r^{p}\right)_{X}.$
We can then further put the first-order system (7.3) into the associated characteristic form which reads
$\frac{\partial r_{1}}{\partial T}+p^{\frac{1}{2}}\overline{r}^{\frac{p-1}{2}}\frac{\partial r_{1}}{\partial X}$ $=0,$ (7.5a)
$\frac{\partial r_{2}}{\partial T}-p^{\frac{1}{2}}\overline{r}^{\frac{p-1}{2}}\frac{\partial r_{2}}{\partial X}$ $=0,$ (7.5b)
@@
$r_{1}$ $=\overline{\rho}-\frac{2p^{\frac{1}{2}}}{p+1}\overline{r}^{\frac{p+1}{2}},$ (7.6a)
$r_{2}$ $=\overline{\rho}+\frac{2p^{\frac{1}{2}}}{p+1}\overline{r}^{\frac{p+1}{2}}.$ (7.6b)
@@
7.3 Initial data for the density variable $\rho$
To set up the initial condition for the variable $\rho$ used in the numerical simulation of the PDE, we have to start from the following crucial jump condition specified in [22]. This jump condition results by demanding that the Riemann invariant $r_{2}$ (with associated characteristic velocity $\lambda_{-}$) of the dispersionless system to be constant. In other words, we must have that
$r_{2}\left(r^{+},\rho^{+}\right)=r_{2}\left(r^{-},\rho^{-}\right),$ (7.7)
where $\rho^{-},\rho^{+}\in\mathbb{R}$ are the constant states in the initial data for the variable $\rho$, namely,
$\rho(X,0)=\begin{cases}\rho^{-}\,,&a<X<b\,,\\ \rho^{+}\,,&X<a~{}\vee~{}X>b\end{cases}$ (7.8)
and $a,b\in\mathbb{R}$ are defined in Eq. (7.1). Notice that the values of $\rho^{-}$ and $\rho^{+}$ are not known at this point, and we have to determine their values according to the jump condition specified in Eq. (7.7). To this end, substituting the expression of $r_{2}$ defined in Eq. (7.6) into the jump condition (7.7), we obtain that,

### v2

@@ 2 Models and theoretical setup
In this work we focus on granular lattice dynamical systems.
The non-dimensional equations of motion (where the constant elastic and geometric prefactors have been absorbed through a suitable rescaling) is given by the following differential-difference equations (DDEs) [5, 11, 26]:
@@
where $\delta_{0}$ denotes the precompression constant, $u_{n}$ the displacement of the $n$th particle from its equilibrium position, and $\left[f\right]_{+}=\text{max}(f,0)$, models the fact that there is no force when the particles come out of contact. The case of $p=3/2$ corresponds to a lattice of spherical particles. It is relevant to mention in passing that in other settings such as O-rings, cylindrical particles or hollow spheres the nature of the force (and hence exponent) may vary [27] and hence we will maintain $p$ as a general parameter in our considerations herein. Equation (2.1) possesses traveling wave solutions [5, 14, 11, 12], even though their exact form is not known analytically (but for relevant approximations see e.g., [28]). Moreover, numerical investigations of Eq. (2.1) reveal the formation of dispersive shocks in certain regimes [7, 8, 9, 26].
In a recent work, we studied the DSWs produced by Eq. (2.1) when $\delta_{0}\neq 0$, making use of two different approximations for it: the Toda lattice and the Korteweg-de Vries equation [25]. In the present work, we focus on the case where $\delta_{0}=0$, in which case the model does not admit a meaningful linear limit in the case of zero background.
It will be convenient to work with the strain variable $r_{n}=u_{n-1}-u_{n}$, which has a physically meaningful interpretation in terms of granular crystals (the amount of compression between adjacent particles), and is common when describing granular crystals, see, e.g., the authoritative book of [5]. We note in passing, however, that some authors also use the variable $u_{n}-u_{n-1}$ for the relevant analysis [18]. This allows one to rewrite Eq. (2.1) as
$\ddot{r}_{n}=\left(r_{n+1}\right)^{p}-2\left(r_{n}\right)^{p}+\left(r_{n-1}\right)^{p},$ (2.2)
where we dropped (here and henceforth) the subscript $+$ under the assumption that the strains are non-negative. The linearized dispersion relation for small-amplitude wave solutions of the form $r_{n}(t)=A+Be^{i(kn-\omega t)}$, with $|B/A|\ll 1$, is given by
@@
$r_{TT}=\left(r^{p}\right)_{XX}+\frac{\epsilon^{2}}{12}\left(r^{p}\right)_{XXXX}=\left(r^{p}\right)_{XX}+\mathcal{O}(\epsilon^{2}),$ (2.10)
@@
$\left(r^{p}\right)_{XXXX}=\left(\left(r^{p}\right)_{XX}\right)_{XX}=\left(r_{TT}-\mathcal{O}(\epsilon^{2})\right)_{XX}=r_{TTXX}+\mathcal{O}(\epsilon^{2}).$ (2.11)
@@ 3.1 Derivation of solitary waves
where denotes differentiation with respect to $Z$ throughout this section, and $a,b$ are two integration constants. When $p$ takes on half-integer values, it is convenient to apply the transformation $R=g^{2}$ so that Eq. (3.2) becomes
@@
where $Z_{0}$ is an integration constant. Indeed, the relevant calculation arises in the process of identifying the bright solitary waves of the general power variant of the nonlinear Schrödinger equation [35]. Recall that since $R=g^{2}$, the corresponding traveling wave solution for $R$ assumes the form
@@ 3.2 Comparison of traveling waves of different models
Figure 2 showcases the numerical comparison of 4 traveling wave solutions (3 approximations and 1 “exact” —again, to a prescribed tolerance—) for the cases of $p=3/2$, $2$, $3$.
@@
Each of the traveling wave solutions which arise in periodic (cosinusoidal) form is plotted only for one period of the corresponding cosine function, beyond which the solution is taken to be zero. Note that, while the continuum models themselves depend on the parameter $\epsilon$, the corresponding predictions is independent of $\epsilon$ once returning to the original lattice variables, see e.g. Eq. (3.7). Thus, we do not need to “select” any particular value of $\epsilon$ when comparing solitary waves.
@@
For the values of $p$ considered here, the approximation from (3.7) is meaningfully proximal to the exact solution, as are the other models. See the black solid curve of Fig. 2. In terms of the error $E=\frac{1}{N}\sum_{n}^{N}|r_{n}-R(\epsilon n)|$, where $r_{n}$ is an “exact” traveling wave and $R$ a traveling wave from one of the PDE models, the approximation errors are of the same magnitude. For example, with $p=3/2$ the errors are $E=0.034$ for the strain variant of the model of Nesterenko (3.11), $E=0.030$ for the model of Ahnert-Pikovsky (2.9), and $E=0.046$ for the model of considered in this paper, namely the regularized model of (2.12). The errors for other values of $p$ are similar. We notice that although the error of the model (2.12) is the greatest, it is still quite small and also comparable to the other two. Recall that the solitary wave solution of the granular chain has a double exponential decay in the tails [37, 38], whereas the approximation from (3.7) has only exponential decay. The approximations (3.9) and (3.13) have finite support. See Fig. 2(d-f), which shows the solutions in a semi-log scale where the decay rate can be better discerned. While the quantitative differences between the actual solution and approximation are similar for all cases, it still remains an interesting open question if, in terms of rigorous error bounds, any of the three considered here is the “most” accurate. It is worth mentioning that here we have only examined the stationary (in the co-traveling frame) aspect of solitary waves. Towards the end of our presentation, we will return to the dynamical properties of these models, as concerns the prototypical structure considered herein, namely the DSW.
@@
Figure 2: Comparison of the traveling solitary waves in different continuum models for different values of the parameter $p$: The leftmost, middle, and rightmost columns denote the cases of $p=3/2,2,3$, respectively. The blue lines (with circles) denote the “exact” solitary of Eq. (3.15), and the red dashed line, green dotted line, black solid line refer to the solitary waves associated with the Ahnert-Pikovsky (2.9), Nesterenko (3.11), and the regularized continuum model (2.12), respectively. The first row shows the comparison of the solitary waves of different continuum models in their respective standard scale, while the second row depicts the semi-log scale of all solitary waves. Note that Ahnert-Pikovsky’s and Nesterenko’s approximations for the solitary waves are only plotted over the period of the respective cosine.
@@ 4.1 Case $p=3/2$
where, as before, primes denote differentiation with respect to $Z$.
A direct integration of Eq. (4.1) yields,
@@
Note that this is only a two-parameter family of periodic waves, and hence is not the most general form. In the derivation of this formula, one of the integration constants was set to zero, see Eq. (4.1). We were unable to find a analytical formula in the general, three parameter, case.
@@ 4.2 Case $p=2$
Note, in contrast to the previous case, this ODE has three free parameter $(B,C,c)$. A direct integration of Eq. (4.4) yields
@@ 4.3 Case $p=3$
Clearly $\mu<0$, and then we first make the assumption that all four roots of $R_{1},R_{2},R_{3},R_{4}$ are real valued and further assume the following order of the four roots,
@@
Here, we need to take the root $R_{4}^{+}$ and ignore $R_{4}^{-}$ to avoid the issue of negative soliton amplitude.
Finally, substituting $R_{4}^{+}$ into Eq. (4.16) we obtain an explicit soliton amplitude formula for $p=3$,
@@ 5 Conservation laws
respectively. We notice that Eq. (5.4) corresponds to the conservation of linear momentum, while Eq. (5.5) refers to the conservation of energy. In addition, it is also worthwhile to note that for the particular case when $\epsilon=0$ and $p=1$, Eq. (5.3) simply reduces to the familiar linear wave equation where this situation immediately falls back to a standard exercise regarding the two conservation laws (see, e.g., [40]).
At the strain level, the PDE (2.12) yields an equivalent conservation law to (5.3):
@@ 6.1 Theoretical preliminaries
The idea of modulation theory is to consider slow modulations of the parameters that completely determine the periodic solutions and derive their associated governing equations. To this end, we first introduce the wavenumber $K$ and frequency $\Omega$ to rewrite the traveling-wave ansatz as $r(X,T)=R(\theta)$, with the phase variable $\theta=(KX-\Omega T)/\epsilon$ where $\Omega=c_{p}K$, to express our periodic solutions.
In the case of $p=2$, the periodic solution in terms of $\theta$ is,
$R\left(\theta\right)=R_{2}+\left(R_{3}-R_{2}\right)\text{cn}^{2}\left(\frac{\sqrt{2\left(R_{3}-R_{1}\right)}}{c_{p}K}\left(\theta-\theta_{0}\right),m\right),$ (6.1)
@@
$2\pi=\int_{0}^{2\pi}\text{d}\theta=2\int_{R_{2}}^{R_{3}}\frac{dR}{R_{\theta}}=2\int_{R_{2}}^{R_{3}}\frac{dR}{\sqrt{\frac{8}{K^{2}{c_{p}}^{2}}\left(R_{1}-R\right)\left(R_{2}-R\right)\left(R_{3}-R\right)}}=\frac{2c_{p}KK_{m}}{\sqrt{2\left(R_{3}-R_{1}\right)}},$ (6.2)
@@
$R\left(\theta\right)=R_{2}+\left(R_{3}-R_{2}\right)\text{cn}^{2}\left(\frac{K_{m}}{\pi}\theta,m\right).$ (6.4)
@@ 6.2 Derivation of modulation equations
To derive the modulation equations, we will use the method of averaged Lagrangian [43]. We first note that the PDE model in its displacement form, see Eq. (5.3), can be obtained through a variational principle. We observe that the Lagrangian density $\mathbb{L}$ associated with Eq. (5.3) is
$\mathbb{L}=\frac{1}{2}\left(u_{T}\right)^{2}+\frac{\epsilon^{2}}{24}u_{XX}u_{TT}-\frac{\left(u_{X}\right)^{p+1}}{p+1}.$ (6.9)
@@
Note that, with this ansatz, all terms can be expressed in terms of the traveling wave at the strain level $R$. In particular, $R=R(\theta)$ satisfies
$K^{2}R_{\theta}^{2}=-\frac{24}{\left(p+1\right)c_{p}^{2}}R^{p+1}+12R^{2}-24BR-12C,$ (6.13)
@@
[…]
$+\left(\frac{c_{p}^{2}}{12}\left(W_{c_{p}}c_{p}+W\right)+\frac{2c_{p}^{2}}{K}\left(-\frac{1}{2}\beta^{2}+\frac{1}{2}C+\beta B\right)\right)_{X}$ $=0,$ (6.25a)
@@ 7.1 Riemann invariants of the dispersionless system
Before discussing numerical simulations and the analytical method of DSW fitting, we need to understand the dispersionless averaged version of the continuum model (2.12). First note that Eq. (2.12) can be written in the following form,
$\left(r-\frac{\epsilon^{2}}{12}r_{XX}\right)_{T}-\rho_{X}$ $=0,$ (7.1)
$\rho_{T}$ $=\left(r^{p}\right)_{X}.$
According to [22], the Whitham modulation equations at both harmonic and solitonic edges should read
@@
These equations are obtained by setting the dispersion term to zero ($\epsilon=0$) in Eq. (7.1) and averaging. We can then further put the first-order system (7.2) into the associated characteristic form which reads
$\frac{\partial q_{1}}{\partial T}+p^{\frac{1}{2}}\overline{r}^{\frac{p-1}{2}}\frac{\partial q_{1}}{\partial X}$ $=0,$ (7.3a)
$\frac{\partial q_{2}}{\partial T}-p^{\frac{1}{2}}\overline{r}^{\frac{p-1}{2}}\frac{\partial q_{2}}{\partial X}$ $=0,$ (7.3b)
@@
$q_{1}$ $=\overline{\rho}-\frac{2p^{\frac{1}{2}}}{p+1}\overline{r}^{\frac{p+1}{2}},$ (7.4a)
$q_{2}$ $=\overline{\rho}+\frac{2p^{\frac{1}{2}}}{p+1}\overline{r}^{\frac{p+1}{2}}.$ (7.4b)
@@
7.2 Initial data for the continuum model
The classic Riemann problem for the continuum model corresponds to the initial conditions
$r(X,0)=\begin{cases}r^{-}\,,&X<0\,,\\ r^{+}\,,&X>0\,,\end{cases}\qquad\rho(X,0)=\begin{cases}\rho^{-}\,,&X<0\,,\\ \rho^{+}\,,&X>0\,,\end{cases}$ (7.5)
These initial conditions are assumed when conducting the DSW fitting analysis in Sec. 7.5. In what follows, we assume that $r_{-}>r_{+}$.
The initial condition for the variable $\rho$ should satisfy a jump condition, as detailed in [22]. In particular, this jump condition is obtained by demanding that the Riemann invariant $q_{2}$ (with associated characteristic velocity $\lambda_{-}$) of the dispersionless system Eq. (7.2) to be constant. In other words, we must have that
$q_{2}\left(r^{+},\rho^{+}\right)=q_{2}\left(r^{-},\rho^{-}\right).$ (7.6)
Notice that the values of $\rho^{-}$ and $\rho^{+}$ are not known at this point, and we have to determine their values according to the jump condition specified in Eq. (7.6). To this end, substituting the expression of $q_{2}$ defined in Eq. (7.4) into the jump condition (7.6), we obtain that,